OpenStudy (anonymous):

The emf in a milli henry inductor in which the current changes from 3A to 1A in a millisecond is:   A) 2 volt         B) 0.2 volt C) 20 volt D) 0.02 volt

OpenStudy (anonymous):

V=-LdI/dt L=1mH, dI/dt = (1-3)/.001 = -2000 A/s V=1mH*2000A/s = 2V